OpenStudy (anonymous):

find where the slop of the curve is defined...: x^2y-xy^2=4

OpenStudy (anonymous):

This requires finding the derivative of y wrt x, i.e. we want to know when dy/dx is definted. We are going to use implicit differentiation: \[ x^2 y - xy^2 = 4 \]\[ 2x \cdot y + x^2 \cdot \dfrac {\mathrm{d}y}{\mathrm{d}x} - x \cdot 2y \cdot \dfrac {\mathrm{d}y}{\mathrm{d}x} - y^2 = 0 \]Can you do the rest?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

No problem! :)